Hi, im Very confused on Linear Graphs anything to do with it.
First here is a problem the i have that i cannot understand at all
{2x+y=-3
{ y= 7
and it says "use substitution to solve the problem"

Found 2 solutions by checkley71, allamericangirl:
Answer by checkley71(8403)

2X+Y=-3 WITH Y=7 THEN SUBSTITUTING 7 FOR Y IN THE FIRST EQUATION WE GET
2X-7=-3 OR 2X=4 OR X=4/2 OR X=2

Answer by allamericangirl(1)

2x+y=-3 for y=7
step 1:substituting 7 for y gives you the equation: 2x+7=-3
step 2:isolate the variable by subtracting 7 from both sides. what you do to one side of an equation you must also do to the other.
step 3: the opposite of the opposite of any integer is the integer itself so we would write out -3-7as -3+(-7) and add.
step 4:divide both sides by 2 to isolate the variable(x)
remember the rule for dividing integers
2x+7=-3
2x+7-7=-3-7
2x=-3+(-7)
2x=-10
2x/2=-10/2
x=-5